OpenStudy (anonymous):

Please help im running out of time! Which equations in point-slope form are equations of the line that passes through the point (4, 5) and (−3, −1)? Choose exactly two answers that are correct. A. y-3=7/6(x-1) B. y-5=6/7(x-4) C. y+1=6/7(x+3) D. y-4=7/6(x-5) I could answer it but this would save me TONS of time, ill answer my other questions while i wait for someone to answer this and i can learn how to do this later.

OpenStudy (anonymous):

Woah instant people

OpenStudy (anonymous):

5 people looked at my post 5 seconds after i posted it up ._.

OpenStudy (haichi):

B

OpenStudy (anonymous):

Any explanation?

OpenStudy (haichi):

yes

OpenStudy (anonymous):

You don't have to, but it would be appreciated c:

OpenStudy (haichi):

point slope form= y-y1=m(x-x1) so I just subistuted

OpenStudy (anonymous):

Oh hm. Thats it?

OpenStudy (haichi):

yeah

OpenStudy (anonymous):

Ok, i understand now. I knew the formula but didnt know it worked with these types of questions too. If anyone else dosent get how to do it, substitute the y1, x1, and the M. (M=slope)

OpenStudy (haichi):

Ye Great job*claps for shawn*

OpenStudy (anonymous):

Oh but wait, this is a multiple choice question...

OpenStudy (anonymous):

Theres supposed to be more than one... Hm...

OpenStudy (haichi):

WEll u can choose either coordinate to be ur x1 and y1

OpenStudy (anonymous):

I guess there dosent have to be more than one, but then why would it make sense to make it multiple choice... Lol, oh well ill try. Yae problemo solved

OpenStudy (anonymous):

Yep, some questions say pick all that are correct and some say pick EXACTLY two. Sorry, i need second awnser. Can anyone help?

OpenStudy (anonymous):

@jim_thompson5910 @dan815 @dumbcow @Compassionate @kropot72 @Kainui @sammixboo

OpenStudy (anonymous):

Kinda panicking ._.

OpenStudy (anonymous):

Ok i think its B and C because when you use both points seperately